OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form.

jimthompson5910 (jim_thompson5910):

slope m = (y2-y1)/(x2-x1) m = (-2-0)/(1-0) m = -2/1 m = -2 ------------------- y = mx+b 0 = -2(0)+b 0 = 0+b 0 = b b = 0 So the equation is y = -2x